Jefferson Lab Weekly Briefs
September 1, 2010
12 GeV UpgradeThe klystron vendor has successfully demonstrated 13 kilowatt output from the initial klystron. This was a significant event for the radiofrequency system’s success. The required operating parameters have been transmitted to the vendor for the power supplies for the klystrons. The first new dipole magnet for the beam transport spreaders and recombiners has been received. Five additional 4-meter beam transport dipole magnets were received, which brings the total delivered to 17 of the 36 in the contract. The 12 GeV accelerator installation work during the current accelerator shutdown is going extremely well. Low-level RF fabrication and Central Helium Liquefier installation activities continue to be ahead of schedule. Testing of the first two niobium cavities for the cryomodule production was suspended while modifications to the Test Lab are being made as part of the TEDF project.
AcceleratorThe second full week of the scheduled down went well. Injector vacuum work was completed, and the system was pumped down with no leaks. The failed chopper housing was removed and replaced. Turbo vacuum carts were fitted with an uninterruptible power supply to allow them to stay on and continue to function during minor power interruptions. Wednesday's planned power outage went smoothly after a rough start. When power was turned off, the backup generator failed to start due to a low coolant level, which was eventually traced to a flaky level sensor. Coolant level was raised to satisfy the sensor and the generator was started within 45 minutes. The 40 MVA substation work was completed and power restored to the normal source by 10:30 a.m. Contractors completed the low-conductivity water piping installation in the west arc. The Machine Control Center UPS was replaced with a new unit. The vacuum valve for 2L16 was replaced and pump downs began in preparation for cooldown next week. Beam switchyard dump shielding was temporarily removed and the dumps were wired to provide a current readback. There were too many cable pulls, alignments and upgrades to list separately, but nonetheless, a lot of hard work is being performed by many.
Free-Electron LaserThe ultraviolet FEL ran extremely well last week. The FEL team lased robustly again and measured the gain at 700 nanometers. The team then shut down for a major sitewide power outage and later brought all systems back successfully. The team took advantage of the shutdown to align the 400 nm mirrors for the push to shorter wavelengths, which was attempted for the first time on Friday.
Theory CenterIn a new paper by Theory Center staff and researchers in Portugal, a relativistic quark model was applied to study electroproduction of the Δ(1600) resonance, which is the first radial excitation of the ground state Δ(1232) in the spin-3/2, isospin-3/2 channel <arXiv:1008.3822 [hep-th]>. The Δ(1600) resonance is revealed through double-pion production in NN collisions and is included in some π N scattering reaction models. In the current model, the electromagnetic N Δ(1600) transition form factors are determined by adding contributions from valence quarks with those from the pion cloud, with the latter estimated by analogy with the Δ(1232) electroproduction reaction.
JLab's Safety Numbers2 Days since Last Recordable Accident (JLab record: 331)
2 Days since Last Lost Workday Accident (JLab record: 676)
JLab Calendar of EventsSept. 6: Labor Day holiday - JLab closed
Sept. 7: Safety Shoe Vendor onsite
Sept. 20: Safety Shoe Vendor onsite
Oct. 5: American Red Cross Blood Drive
Oct. 5: Science Series: Einstein for Everyone
Environment, Safety, Health & QualityVerifying Lab Status During Inclement Weather
In case of inclement weather, you can always check the lab's status through one of the following communication channels:
New Training Record Management System
Accelerator End-of-Run Party